OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f of x equals two divided by x and g of x equals two divided by x.

OpenStudy (zzr0ck3r):

\[f(x) =\frac{2}{x}\\g(x) = \frac{2}{x}\\f(g(x)) = \frac{2}{g(x)}=\frac{2}{\frac{2}{x}}=2*\frac{x}{2}=x\]